OpenStudy (anonymous):

Apebblefallsfromthetopofacliffthatis180mhigh.Thepebble’sheight above the ground is modelled by h(t) 􏰀 􏰂5t2 􏰂 5t 􏰁 180, where h is the height in metres at t seconds since the pebble started to fall. (a) Find the average rate of change of height between 1 s and 4 s. (b) Find h(3). (c) Find the instantaneous rate of change of height at 3 s. (d) Communication: Explain the meaning of each value you calculated for (a), (b), and (c).

sammixboo (sammixboo):

Ah the first part of your problem is jumbled up

OpenStudy (anonymous):

h(t) = -5t2 - 5t + 180,

ganeshie8 (ganeshie8):

Average rate of change between two points is same as the slope of segment joining those points

OpenStudy (anonymous):

How can I do it? Please help me

ganeshie8 (ganeshie8):

Average rate of change between t = 1 and t = 4 : \[\rm \dfrac{h(4) - h(1)}{4-1}\]

ganeshie8 (ganeshie8):

evaluate the given height function at those values and plugin

OpenStudy (anonymous):

Is it -30?

ganeshie8 (ganeshie8):

Correct!

OpenStudy (anonymous):

Wow! Thanks! Is my answer in letter B correct? It's 120 How about letter C?

ganeshie8 (ganeshie8):

120 is right

ganeshie8 (ganeshie8):

are you familiar with derivatives ?

OpenStudy (anonymous):

No

ganeshie8 (ganeshie8):

you need to find derivative of h(t) for instantaneous rate

OpenStudy (anonymous):

How will I do it? Please please help me.

ganeshie8 (ganeshie8):

There are many ways to work part c, im just not sure which way your teacher wants you work it

ganeshie8 (ganeshie8):

this question is from calculus right ?

ganeshie8 (ganeshie8):

i hope its not from precalculus

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

then you should be knowing differentiation

OpenStudy (anonymous):

This is our 2nd lesson I really don't know differentiation

OpenStudy (anonymous):

Ahhh! Is it like the limit as x2 is approaching x1 ?

ganeshie8 (ganeshie8):

you know limits ?

ganeshie8 (ganeshie8):

then for part c, you just need to find below limit : \[\rm \text{instantaneous rate of change at 3s = } \lim\limits_{t\to 3}\dfrac{h(t) - h(3)}{t-3}\]

OpenStudy (anonymous):

Yes

ganeshie8 (ganeshie8):

\[\rm = \lim\limits_{t\to 3}\dfrac{ -5t^2 - 5t + 180 - 120}{t-3}\]

ganeshie8 (ganeshie8):

\[\rm = \lim\limits_{t\to 3}\dfrac{ -5t^2 - 5t + 60}{t-3}\]

ganeshie8 (ganeshie8):

\[\rm = \lim\limits_{t\to 3}\dfrac{ -5(t^2 + t - 12)}{t-3}\]

ganeshie8 (ganeshie8):

\[\rm = \lim\limits_{t\to 3}\dfrac{ -5(t+4)(t-3)}{t-3}\]

ganeshie8 (ganeshie8):

\[\rm = \lim\limits_{t\to 3} -5(t+4)\]

ganeshie8 (ganeshie8):

\[\rm = -5(3+4)\]

ganeshie8 (ganeshie8):

\[\rm = -35\]

ganeshie8 (ganeshie8):

thats the instantaneous rate of change of height at t = 3s

OpenStudy (anonymous):

But can I also use the table ?

OpenStudy (anonymous):

Thankyou again :) How can I explain it on letter D?

ganeshie8 (ganeshie8):

Average rate of change of height is same as `velocity`

ganeshie8 (ganeshie8):

Average rate of change of height betwee two points is same as the `average velocity` between those points

ganeshie8 (ganeshie8):

Since you got -30 for part a, that means the average velocity of the pebble is -30 m/s between 1s and 4s

OpenStudy (anonymous):

How about letters b and d?

OpenStudy (anonymous):

I mean c

ganeshie8 (ganeshie8):

whats your interpretation for part b ?

ganeshie8 (ganeshie8):

how do u interpret h(3) ?

OpenStudy (anonymous):

I think in 3s time the pebble is in the height 120m

ganeshie8 (ganeshie8):

Very good!

ganeshie8 (ganeshie8):

for part c : the velocity of particle at t=3s is -35 m/s

OpenStudy (anonymous):

Yayyyy! I super thank you, :D

ganeshie8 (ganeshie8):

yw:)

OpenStudy (anonymous):

How can I give you a medal? :)

ganeshie8 (ganeshie8):

looksyou already gave me one :) you need to click on "Best response" button next to a reply

OpenStudy (anonymous):

Yah I already did it, Thankyou again :)